OpenStudy (goformit100):

transform to axis inclined at 30 degree, to the original axis. The equation

OpenStudy (goformit100):

\[x^2 + 2 \sqrt{3}xy - y^2 = 2a^2\]

OpenStudy (goformit100):

@amistre64 @ParthKohli @Mertsj @Hero @UnkleRhaukus @experimentX @Callisto @.Sam. @ganeshie8 @Zarkon @JohnnyC123 @mathslover @robtobey @radar @ajprincess @tcarroll010

OpenStudy (anonymous):

Do you have any work done? like show us what you have so we could see what you need help on. :)

OpenStudy (anonymous):

By the way, i thought you said you were going to be offline for a MONTH? O:

OpenStudy (foolaroundmath):

For an axis that is rotated by an angle \(\theta\) (assuming counterclockwise), the new coordinates \((X,Y)\) are given by: \(X = x\cos \theta + y\sin \theta\) \(Y = y\cos \theta - x\sin \theta\)